OpenStudy (anonymous):

Need help on this problem Determine the number(s) between 0 and 2pi where the line tangent to the given function is horizontal f(x)10sinx-10sqrt(3)cosx I know I have to take the derivative of it its I just dont know how I solve it for the horizontal line

OpenStudy (anonymous):

Do you know to calculate the derivative?

OpenStudy (anonymous):

Yes the derivative is \[f'(x)=10\cos(x)+10\sqrt{3}\sin(x)\]

OpenStudy (anonymous):

ok. what is the slope.of horizontal line then?

OpenStudy (anonymous):

Zero

OpenStudy (anonymous):

So do I set x=0?

OpenStudy (anonymous):

well. that's quire typing and i'm on mobile. ill try to help but im slow :(

OpenStudy (anonymous):

Thats ok

OpenStudy (anonymous):

So I have to do this \[0=10\cos(0)-10\sqrt{3}\sin(0)\]

OpenStudy (anonymous):

Correct?

OpenStudy (anonymous):

eh. no. you want to find follow what x makes f'(x) =0

OpenStudy (anonymous):

So then \[0=10\cos(x)+10\sqrt{3}\sin(x)\]

OpenStudy (anonymous):

\[-10\cos(x)=10\sqrt{3}\sin(x)\]

OpenStudy (anonymous):

Then I divide each one by 10 \[-\cos(x)=\sqrt{3}\sin(x)\]

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

so far :) now to the hard part hehe. I hope I would be ble to help this way.. on worst case ill try when i get home

OpenStudy (anonymous):

Ok :D thanks for the guidence

OpenStudy (anonymous):

Ok I got the awnsers {(5/6)pi and (11/3)pi}

OpenStudy (anonymous):

:D awsome thanks a bunch

OpenStudy (anonymous):

cool. :) ye I just realized how blind I got.. \[cos(x) = -\sqrt{3}sin(x) \\ \frac{cos(x)}{sin(x)} = -\sqrt{3} \\ cot(x) = -\sqrt{3} \] damn so slow on mobile -.-